Can someone help me solve this step by step?
Solve the system by the elimination method. First simplify equations, and clear it of fractions.
5x-y=-14+2x+y
7x+9y+4=3x+8y

Answer by Alan3354(69443)

Solve the system by the elimination method. First simplify equations, and clear it of fractions.
5x-y=-14+2x+y
7x+9y+4=3x+8y
--------------
There are no fractions.
You simplify it, then repost it.